A room is 1.5 times as long as it is wide. and it`s perimeter is 25 meters. find the dimension of the room?
Answer by checkley75(3666)

P=2L+2W
25=2(1.5W)+2W
25=3W+2W
25=5W
W=25/5
W=5 IS THE WIDTH
L=1.5W
L=1.5*5
L=7.5 IS THE LENGTH
PROOF
25=2*7.5+2*5
25=15+10
25=25